# Capacitance-based Fermion parity read-out and predicted Rabi oscillations in a Majorana nanowire

### v2

@@ Abstract
Figure 5: Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.3$ for the upper panels, $0.1$ for (c) and $1.1$ for (d)) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. The left panels (right panels) refer to the case for even (odd) total Fermion parity respectively of the system shown in Fig. 1. The top panels (a,b) correspond to parameters of Fig. 3(a,b,c) with inter-wire parity Rabi oscillation dephasing rate $(T_{2}^{\prime})^{-1}\sim 0.02\,GHz$ while the bottom one corresponds to parameters Fig. 3(d,e,f) with $(T_{2}^{\prime})^{-1}\sim 0.42\,GHz$.
@@
Comparing the results for our simulated flux and Fermion parity dependence of the capacitance for the ideal MZM case (i.e. Fig. 2(a,b,c)) to the other cases suggest that the capacitance in the topological case is characterized by an $hc/e-$periodic flux dependence together with an exact $hc/2e$ phase shift between the different parity sectors. Indeed, as shown in the SI Sec. $S_{5}$, both by explicit calculation of low energy ABSs as well as the fractional Josephson effect [9, 1, 2, 3] that this is a definitive signature of MZMs. A gapped trivial phase shows an $hc/2e-$periodic flux dependence similar to Fig. 2(d,e,f). The topological patch regime Fig. 3, while showing a near $hc/2e$ phase shift between the different parity sectors, also shows a change in average as well as amplitude that are not allowed in the topological phase. The recent capacitance measurements in the Majorana nanowire [8], do not distinguish between MZM and the topological patch regime because of an ambiguous flux periodicity.
Furthermore, our analysis establishes that the samples studied in Ref. [8] should manifest Rabi oscillations.In fact, we find in Figs. 35 that the coherence of the Rabi oscillation matches the inter-wire parity coherence $T_{1}^{\prime}$, which in turn is proportional to the gate charge noise. This suggests that in the case of devices containing wires from different categories in Figs. 35 the coherence would be limited by the lower quality nanowire. The combination of flux dependence capacitance, coherent tunneling Rabi oscillation as well as inter-wire parity Rabi oscillations can indicate thus providing a direct qubit diagnostic for Majorana nanowires.
@@ Acknowledgements.
Appendix A Supplementary Information
A.1 $S_{1}$: Details of Majorana nanowire model with quantum well and disorder
@@
The topological stability $T_{S}$ used to characterize the topological nature of the wire is closely related to the topological invariant of the Majorana nanowire defined by Kitaev [1]. The topological invariant or Majorana number was defined by Kitaev for the Hamiltonian in Eq. 3, which has periodic/antiperiodic boundary conditions as the ground state fermion parity of $H$. The ground state fermion parity of $H$ can be written as $P(\lambda_{QD})=\textrm{sign}(Pffafian(\sigma_{y}\tau_{y}H(\lambda_{QD})))$. The topological invariant was defined by Kitaev [1] as the ratio of the Fermion parity between periodic and anti-periodic boundary conditions $P_{top}=P(\lambda_{QD}=1)/P(\lambda_{QD}=-1)$. This topological invariant provides no insight into the proximity to the topological transition. The Pffafian in $P$ changes sign when the smallest eigenvalue of $H(\lambda_{QD}=\pm 1)$ crosses 00. For this reason we define the topological stability
$T_{S}=P_{top}\textrm{min}(E_{g}(\lambda_{QD}=\pm 1))/\textrm{max}(E_{g}(\lambda_{QD}=\pm 1)),$ (4)
where $E_{g}(\lambda_{QD})$ is the gap of the Hamiltonian $H(\lambda_{QD})$.
A.2 $S_{2}$: Technical details of Rabi oscillation calculation
@@
Figure 6: (a) Energy in K and (b) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=500\,nm$ Majorana nanowire in the topological patch regime [11] at $\mu=-1.62\,K$ and $B_{x}=0.28\,T$, $V_{QD}=0.0\,K$, $V_{RMS}=0\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability (defined below Eq. 3) $T_{S}=-0.06$. (c,d) Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.2$ for the upper panels and $14$ for the lower panels) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. Panels (c) and (d) refer to Rabi oscillations in the even and odd total Fermion parity respectively of the system.
A.3 $S_{3}$: Rabi oscillation calibration
@@
A.4 $S_{4}$: Inter-wire parity Rabi oscillations
@@
A.5 $S_{5}$: Characterizing MZM from capacitance
The fractional Josephson effect [9, 1, 2, 3] in the topological superconducting phase can be used to argue for a characteristic flux dependence of the capacitance. To understand this, note that for a ”long” superconducting wire, which is longer than the coherence length, the introduction of a phase slip in the middle of the superconducting wire is equivalent to an $hc/2e$ phase shift. In the trivial phase, this cannot have an effect on the capacitance, which must then be $hc/2e-$ periodic. In contrast, in the topologically non-trivial phase, such a phase slip must be accompanied by a change in fermion parity. Thus, the capacitance in the topological case must be characterized by an $hc/2e-$ phase shift between the two fermion parity sectors, which must otherwise be $hc/e-$periodic.
To check this result explicitly, consider a low energy ABS or pair of MZMs in a wire, which can be represented by a fermion operator $c=(\gamma_{R}+i\gamma_{L})2^{-1/2}$. Let us represent the low energy fermion operator on the site $N=0$ by $d$. Introducing the flux dependent phase $\phi$ (in units so that $hc/2e=2\pi$) on the right tunnel junction:
$H=\delta c^{\dagger}c+2^{1/2}t_{R}e^{i\phi/2}(\gamma_{R}+\zeta_{R}\gamma_{L})d$
$+2^{-1/2}t_{L}d(\gamma_{L}+\zeta_{L}\gamma_{R})+h.c.,$ (11)
$H=\delta c^{\dagger}c+d[t_{R}e^{i\phi/2}\{(1+i\zeta_{R})c^{\dagger}+(1-i\zeta_{R})c\}$
$+t_{L}\{(i+\zeta_{L})c^{\dagger}+(-i+\zeta_{L})c\}]+h.c.$ (12)
Note that choosing $\zeta_{R,L}$ to be complex allows $H$ to contain arbitrary complex normal and Andreev tunneling amplitude at each junction.
The two terms separate naturally by fermion parity sector. In the even fermion parity sector $c^{\dagger}c+d^{\dagger}d=0,2$, the $cd^{\dagger}=c^{\dagger}d=0$ so that writing $d^{\dagger}d=c^{\dagger}c=(1+\sigma_{z})/2$ and $c^{\dagger}d^{\dagger}=\sigma_{+}$ we get
$H_{even}(\phi)=\delta\sigma_{z}/2$
+$+\{t_{R}e^{i\phi/2}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})\}\sigma_{-}+h.c.$ (13)
In the odd fermion parity sector $c^{\dagger}c+d^{\dagger}d=1$, the $c^{\dagger}d^{\dagger}=cd=0$ so that writing $21-d^{\dagger}d=c^{\dagger}c=(1+\sigma_{z})/$ 2 and $dc^{\dagger}=\sigma_{+}$ we get
$H_{odd}(\phi)=\delta\sigma_{z}/2$
+$+\{t_{R}e^{i\phi/2}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})\}\sigma_{+}+h.c.$ (14)
Writing the Hamiltonian in each case as $H=(\delta/2)\sigma_{z}+A\sigma_{+}+A^{*}\sigma_{-}$ we can write the capacitance as
where the amplitude for the even and odd case is
$|A|^{2}$ $=|t_{R}e^{i\phi/2}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})|^{2}\textrm{for even parity}$
$=|t_{R}e^{i\phi/2}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})|^{2}\textrm{ for odd parity}.$ (16)
The $\phi-$dependence of $|A|^{2}$ in both cases can be written in terms of three parameters $P,Q$ and $R$ as $|A|^{2}=P+Q\cos{\phi/2}+R\sin{\phi/2}$ ). In principle, Eqn. 15 together with this parametrization allows a fit for $\delta,P,Q$ and $R$ from the flux dependence of $C(\phi)$ over one period. The parameters $P,Q,R$ for each parity sector combine to form six parameters, which can be combined using Eq. 16 to determine $t_{L,R}$ and the real and imaginary parts of $\zeta_{L,R}$.
Note that the MZM case with $\zeta_{L,R}=0$ is characterized by $P$ and $Q$ remaining unchanged between the two parity sectors while $R$ changes sign. This corresponds to a $2\pi$ shift of $\phi$, which is the essential signature expected for a topological phase.

### v3

@@ III.1 Coherence results
Figure 5: Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.3$ for the upper panels, $0.1$ for (c) and $1.1$ for (d)) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. The left panels (right panels) refer to the case for even (odd) total Fermion parity respectively of the system shown in Fig. 1. The top panels (a,b) correspond to parameters of Fig. 3(a,b,c) with inter-wire parity Rabi oscillation dephasing rate $(T_{1}^{\prime})^{-1}\sim 0.02\,GHz$ while the bottom one corresponds to parameters Fig. 3(d,e,f) with $(T_{1}^{\prime})^{-1}\sim 0.42\,GHz$.
@@
IV Summary and discussion
To summarize our results, motivated by the recent microwave-based Majorana parity read-out experiments [8], we have simulated the results of parity measurement for different phases of the semiconductor nanowire system as well as proposed Rabi oscillations as a path to measuring coherence. We find that while the so-called topological patch regime [11] can produce parity dependent capacitance signals possibly qualitatively similar to experiments, it does not produce an exact $hc/2e$ parity shift that a topological MZMs produce. Improvements to the current parity-read-out experiment [8] that eliminate the background flux dependence as well as being sensitive to the $hc/e$ periodic component of the capacitance signal might help distinguish the MZM from the topological patch regime [8]. The time-averaging in the current experiment [8] leads to the loss of information on the $hc/e$ periodicity so that the claimed periodicity is $hc/2e$. Furthermore, as noted in Sec. II.2, despite similarities to teleportation and fractional Josephson effect, the parity measurement does not directly constrain the MZM splitting $\delta$ beyond temperature. Instead, $\delta$ is determined by transport measurements [25], specifically the zero-bias peak. In this sense, the present capacitance read-out does not contain all features of teleportation proposals [10, 3, 27] where charging energy of the nanowire plays a crucial role.
The single-site approximation discussed in Sec. II, which is valid in the limit of small QD tunneling, was crucial for the systematic simulation using the BdG formalism as discussed in Sec. II. However, it is important to discuss possible limitations of the validity of the conclusions if the tunneling to the QD become large compared to the single-particle level spacing. The approximation focuses on the QD level closest to the Fermi energy and assumes that occupation of other levels can be treated as particle-hole excitations from the QD ground state at a specific number of electrons. Given the energy spacing of the single particle excitations being larger than temperature, one does not expect these excitations to participate as non-virtual excitations at low voltages and frequencies. However, these excitations can become important in a virtual sense in the case where the semiconductor nanowire has a large number of levels in a way similar to the Kondo effect [34], this is not expected to be relevant in this system after it has been studied using transport [25]. Therefore, we do not expect a significant effect of the particle-hole exctiations other than renormalization of the effective tunneling constant that is being used as a tuning parameter to match the magnitude of the capacitance.
The second objective of our work was to study whether the samples studied in Ref. [8] could demonstrate quantum coherence of the parity degree of freedom in the recent experiment. For this we study measurement of inter-wire fermion parity [28] as well as Rabi oscillations using two wires. In fact, preliminary results of the inter-wire parity measurement have already been presented by the Microsoft group [35], though their results are not currently publicly released. While the inter-wire parity measurements are a simpler way to estimate the coherence time that does not rely on coherent inter-wire tunneling, the Rabi oscillations are potentially more distinctive given an oscillatory pattern is simpler to distinguish from noise. One of the conclusions of our Rabi oscillation simulations shown in Figs. 45 is that the coherence of the Rabi oscillation is comparable in magnitude to the inter-wire parity coherence $T_{1}^{\prime}$, which varies by orders of magnitude between the different scenarios considered in Figs. 3 and  5. The inter-wire coherence $T_{1}^{\prime}$, in turn, is proportional to a sum of the intra-wire matrix element of the gate charge noise (see Appendix. D for details) for each wire. This allows us to predict the inter-wire parity coherence time for devices where the two wires are in different phases. Considering Eq. 16 in Appendix D, we conclude that in the case of devices containing wires from different categories in Figs. 45 the dephasing rate can be approximated by the higher decoherence rate wire. Since the Rabi oscillation dephasing rate appears to track the inter-wire decay rate $T_{1}{{}^{\prime}-1}$, we can expect the Rabi oscillation decoherence rate to be dominated by the poorer quality wire in the device as well.
The combination of flux dependence capacitance, coherent tunneling Rabi oscillation as well as inter-wire parity Rabi oscillations can indicate thus providing a direct qubit diagnostic for Majorana nanowires. The signature of MZM then becomes an operational one i.e. enhancement of coherence by topology. A strong topological phase should manifest as an enhanced robustness to charge noise. Such an observation of orders of magnitude enhancement of the coherence time as calculated in our ideal MZM example (Fig. 4 would be a very strong indication of a topological superconducting phase.
@@ Acknowledgements.
Appendix A Details of Majorana nanowire model with quantum well and disorder
@@
The topological stability $T_{S}$ used to characterize the topological nature of the wire is closely related to the topological invariant of the Majorana nanowire defined by Kitaev [1]. The topological invariant or Majorana number was defined by Kitaev for the Hamiltonian in Eq. 9, which has periodic/antiperiodic boundary conditions as the ground state fermion parity of $H(\phi)$. The ground state fermion parity of $H$ can be written as $P(\phi)=\textrm{sign}(\textrm{Pffafian}(\sigma_{y}\tau_{y}H(\phi)))$. Since, shifting $\phi$ by 1, the topological invariant defined by Kitaev [1] can be written as the ratio of the Fermion parity between periodic and anti-periodic boundary conditions $P_{top}=P(\phi)/P(\phi+1)$. This topological invariant provides no insight into the proximity to the topological transition. The Pffafian in $P$ changes sign when the smallest eigenvalue of $H(\phi=0,1)$ crosses 00. For this reason we define the topological stability
$T_{S}=P_{top}\textrm{min}(E_{g}(\phi=0,1))/\textrm{max}(E_{g}(\phi=0,1)),$ (10)
where $E_{g}(\phi)$ is the gap of the Hamiltonian $H(\phi)$.
Appendix B Technical details of Rabi oscillation calculation
@@
Figure 6: (a) Energy in K and (b) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=500\,nm$ non-superconducting (i.e. $\Delta_{0}=0$) nanowire in the topological patch regime [11] at $\mu=-1.62\,K$ and $B_{x}=0.28\,T$, $V_{QD}=0.0\,K$, $V_{RMS}=0\,K$. The wave-functions for this case shows a single low-energy bound state with a topological stability (defined below Eq. 9) $T_{S}=-0.06$. (c,d) Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.2$ for the upper panels and $14$ for the lower panels) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. Panels (c) and (d) refer to Rabi oscillations in the even and odd total Fermion parity respectively of the system.
Appendix C Rabi oscillation calibration
@@
Appendix D Inter-wire parity Rabi oscillations